OpenStudy (anonymous):

Factor the polynomial completely. z^3(y^2 − 9) + 125(y^2 − 9)

OpenStudy (anonymous):

y^2-9=(y+2)(y-2) z^3+125=(z+5)^3 (y+2)(y-2)(z+5)^3

OpenStudy (anonymous):

\[(y+2)(y-2)(z+5)^{3}\]

jhonyy9 (jhonyy9):

\[(z ^{3}+125)(y ^{2}-9)=(z+5)(z ^{2}-5z+25)(y-3)(y+3)\]